The population mean and standard deviation are given below. Find the required probability and determine whether the given sample mean would be considered unusual. For a sample of n equals 62​, find the probability of a sample mean being less than 22.9 if mu equals 23 and sigma equals 1.33. For a sample of n equals 62​, the probability of a sample mean being less than 22.9 if mu equals 23 and sigma equals 1.33 is
Answer #1

Solution :

Given that ,

mean = \mu = 23

standard deviation = \sigma = 1.33

n = 62

\mu\bar x = 23

\sigma\bar x =\sigma  / \sqrt n = 1.33/ \sqrt 62=0.17

P(\bar x < 22.9) = P[(\bar x - \mu \bar x ) / \sigma \bar x < (22.9-23) / 0.17]

= P(z < -0.59)

Using z table

= 0.2776